OpenStudy (anonymous):

A particles velocity is given by the function : v=kt^2 (v= k(t-squared) where k is a constant and t is in seconds. The particle's position at t=0, is -9 m. At t =3, the position is 9m. Determine the value of the constant k, and be sure to include the proper units.

OpenStudy (fifciol):

start from this equation : \[v=\frac{ dx }{ dt }\Rightarrow vdt=dx\Rightarrow \int vdt=\int dx \Rightarrow x=\int vdt\]

OpenStudy (fifciol):

substitute v=kt^2 and solve that integral \[x=\int\limits kt^2dt=\frac{ 1 }{ 3 }kt^3+C\]

OpenStudy (fifciol):

then substitute : -9= 1/3 *k*0^3 +C 9=1/3*k*3^3+C so from 1-st equation you can immidiately notice that: C=-9 and by substituting that to 2-nd equation you'll get k=2

OpenStudy (fifciol):

since v=kt^2 k=v/t^2 so the unit of k is m/s/s^2=m/s^3

OpenStudy (anonymous):

thanx @fifciol, sorry I wasn't online for a step by step tutorial